# MotionMap: Representing Multimodality in Human Pose Forecasting

### v1

@@ 4.4.2 Training
Training the heatmap predictor $\mathcal{H}$ to obtain MotionMaps for different samples is depicted in Figure 2. For a given sample $X$, the MotionMap model $\mathcal{H}$ is trained to minimize the pixel wise binary cross entropy loss with the heatmap constructed using ground-truth multimodal samples. To prevent overfitting, the MotionMap model uses a simple GRU encoder which spatio-temporally encodes the last three frames of the observation $X$. This is followed by a single fully connected layer, which culminates in a series of 1x1 convolutions. The architecture is detailed in the supplementary material.
@@ 4.5 Inference
At test time, we use the predicted MotionMap coupled with the codebook to obtain the missing latent $z_{y}$. Specifically, given an observation $X$, we predict the corresponding MotionMap using the model $\mathcal{H}$. We then compute various local maxima deterministically which correspond to the likeliest modes of the future sequences. The number of maxima can vary for different observed sequences. Next, we use the codebook $B$ to index a latent vector for the given maxima, giving us the missing latent $z_{y}$. We also obtain $z_{x}$ from $\mathcal{E}_{X}$, allowing us to pass the updated latent $f(z_{x}\bigoplus z_{y})$ to the decoder for pose forecasting. Finally, we note that this methodology does not rely on stochastic models such as diffusion.
@@ 5 Experiments
The goal of this study is to present human pose forecasting as a well posed problem, which in our scenario is to learn and translate different unseen transitions from the observed dataset to unseen test samples. To this effect, we compare our method against multiple state-of-the-art baselines across the AMASS and the Human3.6M datasets.
  • Human3.6M [33] consists of motion-captured poses of seven publicly available subjects performing 15 actions. We follow the protocol proposed by [1]. The first five publicly available subjects (S1, S5, S6, S7, S8) of the dataset are used for training and the last two (S9, S11) for testing. The dataset consists of 32 keypoints in total, from which 17 are selected. We zero-center them around the pelvis joint, and thus the the remaining 16 joints are forecasted with respect to the pelvis. Videos of this dataset have been recorded at 50 fps, and we take 0.5 seconds (25 frames) as input and forecast the next 2 seconds (100 frames).
5.1 Metrics
We evaluate our results using metrics introduced in prior work [3]. The Average and Final Displacement Error (ADE / FDE) measure the minimum L2 distance among the $K$ predicted pose sequences across all frames and the last frame, respectively. However, this paper focuses on the multimodal versions of these metrics called MMADE and MMFDE, which measure the average ADE and FDE across different multimodal ground truths. A prediction will be penalized if it is not close to any of the test ground truths. We use the standard threshold of 0.5 for Human3.6M and 0.4 for AMASS multimodal ground truth selection.
5.2 Implementation Details
We implemented our methods with PyTorch and ran our experiments on an A100 GPU. The encoding-decoding module was trained for 100 epochs for Human3.6M and 50 epochs for AMASS, with a batch size of 32. We used a learning rate starting from 0.001 and decreasing following a decay on the plateau learning rate schedule. The encoder and decoder are single-layer Gated Recurrent blocks based on [1] and modified for our method. The Heatmap model which predicts the MotionMap is a combination of two simple MLP layers which project the input to the dimensionality of the heatmap, followed by five simple 1x1 convolutional layers.
Figure 7: Diversity. By virtue of using the same decoder as BeLFusion [1], our method learns realistic yet diverse motions. This is because MotionMap can be decoded to select modes which are sufficiently different yet likely.
@@ 5.3 Controllability
With Figure 4, we show that the set of pose forecast predictions can be controlled based on user preference. For example, meta data such as action labels can be used to select from within the likeliest pose forecasts based on user preference. Additionally, forecasts similar to a selected one can be sampled by picking the latents corresponding to the nearest neighbors of the local maxima.
@@ 5.4 A Tale Of Two Uncertainties
With the proposed approach, we not only learn the aleatoric uncertainty for each prediction but also the confidence of the mode itself. Specifically, suppose $Y$ is a predicted motion, and $X$ is the observed ground truth. In that case, we model the variables as $X\rightarrow Z\rightarrow Y$, where $Z$ is the variable that represents the observed mode. Introducing an intermediate variable allows us to convert a multimodal distribution over $Y$ to an approximate unimodal distribution Figure 6. We observe this difference in modeling by comparing [16] with our approach. In the absence of any additional conditioning information, the predicted variance [16] increases till it saturates at the end of the prediction horizon. This arises due to increased ambiguity in the motion as time increases. In comparison, by conditioning on the mode, we limit the set of future motions, allowing for better modelling of the uncertainty with time. Moreover, a heatmap represents a spatial distribution over different modes, allowing us to quantify the likelihood of the mode itself. The increased uncertainty during the transition from observation (time frames 25-30) to prediction allows the model to smoothly transition and avoid discontinuities in motion. The uncertainty is observed to change smoothly when the future selected poses are the actual ground truth of the input sequence. It is also observed that the uncertainty in the joints closer to the pelvis is generally lower due to limited mobility. However, joints with a higher degree of mobility tend to have higher uncertainties. We provide more visual trends of uncertainty for more sequences and per all joints in the supplementary material.
@@ 5.5 Ranking Predictions and Diversity
Unlike state-of-the-art methods, MotionMap by virtue of predicting the confidence of different modes allows us to sort our predictions in order of confidence. We denote this in Figure 5, where we plot the predicted MotionMap along with the associated pose forecasts, ranked based on the confidence of the mode. We observed that high confidence modes are often correlated with the ground truth multimodal future, with very low confidence modes at times resulting in modes which have a rare occurence.
Table 1: Human3.6M dataset: All baselines (except for zero-velocity) are limited to 7 forecasts. Our method, unconstrained by the number of modes, is adjusted to produce an equal number of predictions. Metrics are reported in meters.
Method Diversity ($\uparrow$) ADE ($\downarrow$) FDE ($\downarrow$) MMADE ($\downarrow$) MMFDE ($\downarrow$)
Zero-Velocity 0.000 0.597 0.884 0.617 0.879
TPK [35] 6.533 0.534 0.691 0.559 0.675
DLow [3] 11.770 0.445 0.730 0.576 0.715
GSPS [36] 14.97 0.512 0.684 0.550 0.665
DivSamp [4] 15.733 0.480 0.685 0.542 0.671
BeLFusion [1] 7.107 0.441 0.597 0.491 0.586
CoMusion [2] 7.325 0.426 0.613 0.531 0.623
MotionMap 7.965 0.472 0.594 0.464 0.529
Table 2: AMASS dataset: All baselines (except for zero-velocity) are limited to 7 forecasts. Our method, unconstrained by the number of modes, is adjusted to produce an equal number of predictions. Metrics are reported in meters.
Method Diversity ($\uparrow$) ADE ($\downarrow$) FDE ($\downarrow$) MMADE ($\downarrow$) MMFDE ($\downarrow$)
Zero-Velocity 0.000 0.755 0.992 0.778 0.996
TPK [35] 8.570 0.519 0.634 0.600 0.678
DLow [3] 12.694 0.471 0.594 0.554 0.633
GSPS [36] 13.550 0.501 0.662 0.591 0.688
DivSamp [4] 25.901 0.479 0.638 0.623 0.728
BeLFusion [1] 7.917 0.347 0.478 0.488 0.564
CoMusion [2] 7.390 0.311 0.460 0.526 0.602
MotionMap 8.821 0.324 0.447 0.448 0.510
We have visualized some examples of the generated poses in Figure 7, given test input sequences for qualitative comparison of the generated pose sequences. It can be seen that our generated pose sequences are markedly diverse while also depicting smooth and realistic motions.
@@ 5.6 Quantitative Results
We quantitatively evaluate different baselines on their ability to translate multimodality from the observed data for any unseen test sample. We implement this by calculating multimodal ground truths closest to the testing sample from within the training labels. Such an evaluation not only makes the problem well posed, but also addresses acute mismatches between the training and testing distributions, which we show in the supplementary material (Fig. 13). Moreover, in such scenarios, since the testing multimodal ground truth is not representative of the training data, the testing multimodal ground truth does not encompass a majority of the modes. Let us assume that the testing split has only 5 samples. Then, any sample in the testing data will have at most five and a minimum of one multimodal ground truth.
We compare various baselines in Table 2 and Table 1, corresponding to the Human3.6M and AMASS datasets, respectively. We make two observations: While methods such as DLow and DivSamp are diverse, they do not accurately predict the futures corresponding to the observed motion, since not all anchors in the latent are equally likely. In contrast, by imposing a ’prior’ on the latent, MotionMap successfully predicts the likeliest modes while being sample efficient. We additionally note that MotionMap successfully outperforms strong diffusion-based baselines in making correct multimodal transitions learnt from the observed data over all test samples. Finally, we also report the performance using the existing methodology of restricting the multimodal ground truth to the testing split (Tables 3, 4). We observe that while MotionMap is much more in recalling transitions from the test set, this does not come at the cost of general performance for unseen samples.
Limitations. A primary limitation of our method is the lack of fine-grained motion prediction within a mode, such as predicting subtle variations in walking. We particularly observe this effect in the testing split with many of the multimodal ground truths being time-shifted versions of the observation. These ground truths are clumped by MotionMap under one mode. Since MotionMap encodes only one forecast per mode, this may result in higher “errors” since the forecast cannot explain all subtle variations with the samples. However, the choice of clumping similar motions together remains a design choice of MotionMap.
@@ 6 Conclusion
In this work, we discussed a paradigm to make the learning and evaluation of human pose forecasting well-posed. Next, we proposed a new representation that successfully encodes various future motions per sample. In addition, the representation allows for quantifying the confidence of different modes, as well as successfully learning diverse transitions between the input and observed pose sequence present in the observed data. Finally, by explicitly predicting the spread of future motions, MotionMap is sample-efficient and does not rely on repeated random sampling to achieve mode coverage. Our experiments showed that MotionMap successfully learns to generalize transitions between pose sequences in the observed data and provides suitable measures of uncertainty and confidence per mode.
@@ Appendix A Algorithm: Motion Transfer
/* Transfers the motion from pose sequence $y$ to skeleton of sequence $x$ */
0.2cm // Pose sequence for skeletal reference
@@
$\rho$, _\_, _\_ = cartesian_to_spherical(pose)
@@
_\_, $\theta$, $\phi$ = cartesian_to_spherical(y)
@@ Appendix B Implementation Details
Appendix C Additional Quantitative Results
We report our results by restricting the multimodal ground truth to the testing split only. We observe that across both datasets the quantitative results are similar across different methods. While DivSamp is highly diverse, this does not necessarily translate to accurately predicting possible futures. A major observation is that while MotionMap is much more effective in recalling transitions from the test set (Tables 2, 1), this does not come at the cost of general performance, as evident by these results (3. 4). Finally, we note that restricting the multimodal ground truth to the testing split limits the diversity of modes in the ground truth. In Figure 13, we demonstrate that the AMASS testing dataset does not adequately represent the training data, with the testing multimodal ground truth missing the majority of modes. Assuming that the test split contains only five samples, each test sample would have between one and five multimodal ground truths. Furthermore, a discrepancy in the distributions of the train and test split means that the multimodal ground truths for the test set share no commonalities with the train set.
@@ Appendix C Additional Quantitative Results
Appendix D Additional Qualitative Results
@@ D.1 Controllability
Figure 8: Red crosses mark the modes selected by the model. By hovering over the demo tool, we can view the decoded future poses corresponding to the given input pose sequence. We have uniformly selected eight frames in each sequence to demonstrate the motion and stacked them on top of each other at the end(the frame on the very right of each visualized sequence) to represent the amount of motion in each sequence.
Figure 9: We show different strategies for controlled selection of the other forecasts: (a) Selecting samples in the vicinity of a model selected mode. (b,c,d) Based on the distribution of action labels. For instance, we could generate futures for rarer transitions such as sitting down (b) on a chair (c) on the floor or (d) lying on the floor.
D.2 Uncertainty
@@
D.3 Heatmap Comparison
We compare the predicted MotionMaps with the ground truth heatmaps in Figure 10. MotionMap is encouraged to predict a higher number of modes than present in the ground truth to identify rare transitions. Our visualizations confirms that MotionMap identifies other transitions while not missing out on the original ground truth motions. These miscellaneous transitions are learn by the MotionMap model from trends across the dataset.
Figure 10: Qualitative comparison between the MotionMap predicted heatmap and the ground truth multimodal heatmap. Our observations indicate that MotionMap effectively captures the diversity of the modeled scenarios. The presence of a larger number of peaks in MotionMap corresponds to the larger diversity of multimodal ground truth found in the train split.
How well can state-of-the-art baselines predict multimodality without explicitly encoding multimodal transitions? To study this, we collected 50 predictions for each of the baselines for each input pose seqeuence. We then encode these pose forecasts into two dimensions as described in Section 4.4.1. Next, we overlay them on the ground truth heatmap to identify the differences in the predictions and the ground truth. We observe that baselines that rely on anchors although diverse predict transitions which are unlikely for the given pose sequence. This also tallies with our quantitative evaluation. While this effect is reduced for diffusion based baselines, the methods are less diverse and do not capture rare modes. In contrast, MotionMap captures both common and rare mode since they are encoded in the form of local maxima.
Figure 11: We overlay 50 predictions for each baseline on the ground truth heatmap, for each of the three input pose sequences. The encoding of these 50 predictions is marked with red crosses. For MotionMap, we directly overlay the predicted MotionMap (with crosses for maxima) on the ground truth heatmap. It is observed that methods are either highly diverse but predict unrealistic forecasts or are less diverse but predict likely futures. In contrast, MotionMap predicts both: common and rare modes since both are explictly encoded in the training process.

### v2

@@ 4.2.2 Training
Training the heatmap predictor $\mathcal{H}$ to obtain MotionMaps for different samples is depicted in Figure 2. For a given sample $X$, the MotionMap model $\mathcal{H}$ is trained to minimize the pixel-wise binary cross-entropy loss with the heatmap constructed using ground-truth multimodal samples. To prevent overfitting, the MotionMap model uses a simple GRU encoder which spatio-temporally encodes the last three frames of the observation $X$. This is followed by a single fully connected layer, which culminates in a series of 1x1 convolutions. The architecture is detailed in the appendix.
4.3 Fine-Tuning
To improve the performance at test time, we recommend fine-tuning $f(\cdot)$ and $\mathcal{D}$. This is because the codebook maps the heatmap to the average embedding, =$\mathcal{B}=\{h_{y}:\overline{z_{y}}\}$. The fine-tuning is similar to training, with the exception that we use $\overline{z_{y}}$ instead of $z_{y}$. This requires mapping a multimodal future to its heatmap embedding, and using the codebook to decode the average latent.
@@ 4.4 Inference
At test time, we use the predicted MotionMap coupled with the codebook to obtain the missing latent $\overline{z_{y}}$. Specifically, given an observation $X$, we predict the corresponding MotionMap using the model $\mathcal{H}$. We then compute various local maxima deterministically which correspond to the likeliest modes of the future sequences. The number of maxima $\widehat{M}$ can not only vary for different observed sequences but may also differ from the number of ground truth modes $M$. This is because we use weighted binary cross-entropy to penalize false negatives more than false positives. This leads to potential new modes that emerge across the dataset.
Next, for each mode, we use the codebook $B$ to index the missing latent $\overline{z_{y}}$. We also obtain $z_{x}$ from $\mathcal{E}_{X}$, allowing us to pass the updated latent $f(z_{x}\bigoplus\overline{z_{y}})$ to the decoder for pose forecasting. Finally, we note that this methodology does not depend on stochasticity.
@@ 5 Experiments
Overview. We study [^1] https://www.github.com/vita-epfl/MotionMap human pose forecasting as a well posed problem, which is to learn and translate different unseen transitions from the observed dataset to unseen test samples. We perform extensive quantitative and qualitative analysis related to controllability, ranking, and uncertainty estimation. We use the AMASS  [38] and the Human3.6M  [39] datasets. We describe these implementation and dataset in detail in Appendix B and C. Our evaluation uses metrics from prior work  [3]. Average Displacement Error (ADE) measures the minimum L2 distance between ground truth and predicted sequences averaged over all frames, while Final Displacement Error (FDE) considers only the last frame. This paper specifically emphasizes their multimodal counterparts, MMADE and MMFDE, which average ADE and FDE across multiple ground truths. We apply standard thresholds of 0.5 for Human3.6M and 0.4 for AMASS multimodal ground truth selection.
Figure 7: Diversity. By virtue of using the same decoder as BeLFusion [1], our method learns realistic yet diverse motions. This is because MotionMap can be decoded to select modes that are sufficiently different yet likely.
@@ 5.1 Controllability
Figure 4 demonstrates that pose forecast predictions can be done based on user preferences. For instance, metadata such as action labels can be used to select the most likely pose forecasts that align with user intent. These preferences may stem from desired actions extracted via language models to generate specific pose forecasts. Additionally, forecasts similar to a selected one can be obtained by choosing the latent embedding corresponding to the nearest neighbors of a local maximum. Additional figures are in Appendix E.1
@@ 5.2 A Tale Of Two Uncertainties
In our approach, we learn both the uncertainty of each prediction and the confidence associated with each mode. Specifically, we model the relationship as $X\rightarrow Z\rightarrow Y$, where $Z$ represents the mode and $Y$ represents the forecast. Introducing $Z$ simplifies a multimodal distribution over possible futures into an approximately unimodal distribution. This improves upon homoscedastic modeling, which treats uncertainty as independent of the observed pose [16] and averages uncertainty across all future motions, leading to generic trends such as increasing uncertainty over time.
In contrast, MotionMap captures both action ambiguity and motion-specific uncertainty. It models a spatial distribution over possible actions that also gives the confidence per mode. We also condition on specific modes to limit future motions, resulting in uncertainty trends tailored to the prediction. For instance, the nose joint primarily provides skeleton orientation. In Figure 6, Pred. 4 exhibits greater nose uncertainty than Pred. 6, due to a sharper turn. Additionally, increased uncertainty during the transition (frames 25-30) prevents motion discontinuities between the observed sequence and different multimodal ground truths across various subjects. We also note that joints with higher mobility generally display greater uncertainty. Additional visual examples are provided in Appendix E.2.
5.3 Sampling Comparison
Can we use the MotionMap representation to compare multimodality across state-of-the-art? To do so, we collected predictions for each baseline for three different sequences. We then encode these predictions into two dimensions as described in Section 4.2.1. Next, we overlay them on the ground truth MotionMap to identify the differences in the predictions and the ground truth. We observe in Figure 8 that baselines that rely on anchors, although diverse, predict transitions that are unlikely for the given pose sequence. Diffusion based methods are less diverse and do not capture rare modes. In contrast, MotionMap predictions captures the spread of the multimodal ground truth. The full comparison is in Appendix E.3.
@@ 5.4 Ranking Predictions and Diversity
Unlike state-of-the-art methods, MotionMap predicts the confidence of different modes and allows us to rank our predictions in order of confidence. In Figure 5, we plot the predicted MotionMap along with the associated pose forecasts, ranked based on the confidence of the mode. We observed that high confidence modes are often correlated with the ground truth future, and low confidence modes indicating rare transitions from the observed pose. We also visualize some examples of the generated poses in Figure 7. Given input sequences, we see that our generated pose sequences are markedly diverse while also depicting smooth and realistic motions.
Table 1: Human3.6M dataset and AMASS dataset. All baselines (except for zero-velocity) are limited to 7 forecasts. Our method, unconstrained by the number of modes, is adjusted to produce an equal number of predictions. Metrics are reported in meters.
Human3.6M AMASS
Method Diversity $\uparrow$ ADE $\downarrow$ FDE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ Diversity $\uparrow$ ADE $\downarrow$ FDE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$
Zero-Velocity - 0.597 0.884 0.617 0.879 - 0.755 0.992 0.778 0.996
TPK [40] 6.53 0.534 0.691 0.559 0.675 8.57 0.519 0.634 0.600 0.678
DLow [3] 11.77 0.445 0.730 0.576 0.715 12.69 0.471 0.594 0.554 0.633
GSPS [41] 14.97 0.512 0.684 0.550 0.665 13.55 0.501 0.662 0.591 0.688
DivSamp [4] 15.73 0.480 0.685 0.542 0.671 25.90 0.479 0.638 0.623 0.728
STARS [28] 15.85 0.508 0.697 0.551 0.686 - - - - -
TCD [27] 7.63 0.603 0.828 0.668 0.835 - - - - -
BeLFusion [1] 7.11 0.441 0.597 0.491 0.586 7.92 0.347 0.478 0.488 0.564
CoMusion [2] 7.32 0.426 0.613 0.531 0.623 7.39 0.311 0.460 0.526 0.602
MotionMap 7.84 0.474 0.598 0.466 0.532 8.50 0.325 0.450 0.450 0.514
Figure 8: Sampling comparison. We compare the pose forecasts of different methods using the MotionMap representation for three different observations (a, b, c). The multimodal ground truth is represented through MotionMap, and the red points are two-dimensional projections of the predicted futures. For the MotionMap method, we overlay the predicted MotionMap itself. We observe that the MotionMap model performs better since the predicted MotionMap aligns with the ground truth. Please zoom!
@@ 5.5 Quantitative Results
We quantitatively assess different baselines on their ability to translate multimodality from observed data to any unseen test sample. To achieve this, we identify multimodal ground truths in the training labels that are closest to each test sample. This approach not only ensures a well-posed problem but also mitigates discrepancies between training and testing distributions, as demonstrated in the appendix (Figure 14). Moreover, in such cases, the limited size of the test set may fail to capture the full diversity of modes present in the training data. For instance, if the test split contains only five samples, each test sample will have at most five and at least one multimodal ground truth, making the evaluation inherently unreliable.
We compare various baselines in Table 1 using the Human3.6M and AMASS datasets. While methods such as DLow and DivSamp are diverse, they do not accurately predict the futures corresponding to the observed motion, since not all anchors in the latent are equally likely. These methods also tend to be more diverse than the ground truth itself. In contrast, by imposing a ‘prior’ on the latent, MotionMap successfully predicts the likeliest modes while being sample efficient. We additionally note that MotionMap successfully outperforms strong baselines in extending multimodal transitions in the observed data to all test samples. We also highlight a comparison of MotionMap with BeLFusion [1]. Since the two share the same encoder and decoder, the major difference is in generating $z_{y}$. BeLFusion uses diffusion and repeated sampling whereas MotionMap uses heatmaps and codebooks, providing significant gains.
Finally, we also report the performance using the existing methodology of restricting the multimodal ground truth to the testing split (Tables 2, 3). We observe that while MotionMap is more accurate in recalling transitions from the test set, this does not come at the cost of general performance for unseen motions.
Limitations. A primary limitation of our method is the lack of fine-grained motion prediction within a mode, such as predicting subtle variations in walking. We particularly observe this effect in the testing split with many of the multimodal ground truths being time-shifted versions of the observation. These ground truths are clumped by MotionMap under one mode. Since MotionMap encodes only one forecast per mode, this may result in higher “errors" since the forecast cannot explain all subtle variations with the samples. However, the choice of clumping similar motions together remains a design choice of MotionMap.
@@ 6 Conclusion
In this work, we discussed a paradigm to make human pose forecasting well-posed. Next, we proposed a new representation that explicitly encodes a varying number of future motions for different observations. The representation also quantifies the confidence of different modes, allowing us to identify more likely futures from less likely ones. With MotionMap, we learned diverse yet realistic motion transitions from the observed data. By explicitly predicting the spread of future motions, MotionMap is sample-efficient and does not rely on repeated random sampling to achieve mode coverage. We highlighted a potential use of MotionMap for controllable pose forecasting, as well as uncertainty estimation from a practical standpoint. Through qualitative and quantitative analysis, we highlighted the effectiveness of MotionMap for robust and safe human pose forecasting.
@@ Acknowledgment
Appendix
We make the code available on https://github.com/vita-epfl/MotionMap.
@@ Appendix A Algorithm: Motion Transfer
/* Transfers the motion from pose sequence $y$ to skeleton of sequence $x$/
// Pose sequence for skeletal reference
@@
$\rho$, __\__, __\__ = cartesian_to_spherical(pose)
@@
__\__, $\theta$, $\phi$ = cartesian_to_spherical(y)
@@ Appendix B Implementation Details
Appendix C Dataset: Details
Human3.6M [39] consists of motion-captured poses of seven publicly available subjects performing 15 actions. We follow the protocol proposed by [1]. The first five publicly available subjects (S1, S5, S6, S7, S8) of the dataset are used for training, and the last two (S9, S11) for testing. The dataset consists of 32 keypoints in total, from which 17 are selected. We zero-center them around the pelvis joint, and thus the the remaining 16 joints are forecasted with respect to the pelvis. Videos of this dataset have been recorded at 50 fps, and we take 0.5 seconds (25 frames) as input and forecast the next 2 seconds (100 frames).
Appendix D Additional Quantitative Results
We report our results by restricting the multimodal ground truth to the testing split only. We observe that across both datasets the quantitative results are similar across different methods. While DivSamp is highly diverse, this does not necessarily translate to accurately predicting possible futures. A major observation is that while MotionMap is much more effective in recalling transitions from the test set (Table 1), this does not come at the cost of general performance, as evident by these results (2. 3). Finally, we note that restricting the multimodal ground truth to the testing split limits the diversity of modes in the ground truth. In Figure 14, we demonstrate that the AMASS testing dataset does not adequately represent the training data, with the testing multimodal ground truth missing the majority of modes. Assuming that the test split contains only five samples, each test sample would have between one and five multimodal ground truths. Furthermore, a discrepancy in the distributions of the train and test split means that the multimodal ground truths for the test set share no commonalities with the train set.
@@ Appendix D Additional Quantitative Results
Appendix E Additional Qualitative Results
@@ E.1 Controllability
Figure 9: We visualize the modes (in red crosses) predicted by MotionMap. By hovering over the demo tool, we can view the decoded future poses corresponding to the given input pose sequence. We have uniformly selected eight frames in each sequence to demonstrate the motion and stacked them on top of each other at the end (the frame on the very right of each visualized sequence) to represent the amount of motion in each sequence.
Figure 10: We show different strategies for controlled selection of non-maxima forecasts: (a) Selecting samples in the vicinity of a model selected mode. (b,c,d) Based on the distribution of action labels. For instance, we could generate futures for rarer transitions such as sitting down (b) on a chair (c) on the floor, or (d) lying on the floor.
E.2 Uncertainty
@@
E.3 Sampling Comparison
We compare the predicted MotionMaps with the ground truth heatmaps in Figure 11. MotionMap is encouraged to predict a higher number of modes than present in the ground truth to identify rare transitions. Our visualizations confirms that MotionMap identifies other transitions while not missing out on the original ground truth motions. These miscellaneous transitions are learned by the MotionMap model from trends across the dataset.
Figure 11: Qualitative comparison between MotionMap and the ground truth multimodal heatmap. Our observations indicate that MotionMap effectively captures the diversity of the modeled scenarios. The presence of a larger number of peaks in MotionMap is a result of learning generalized behaviour across the training split.
How well can state-of-the-art baselines predict multimodality without explicitly encoding multimodal transitions? To study this, we collected predictions for each of the baselines for each input pose seqeuence. We then encode these pose forecasts into two dimensions as described in Section 4.2.1. Next, we overlay them on the ground truth heatmap to identify the differences in the predictions and the ground truth. We observe that baselines that rely on anchors although diverse predict transitions which are unlikely for the given pose sequence. This also tallies with our quantitative evaluation. While this effect is reduced for diffusion based baselines, the methods are less diverse and do not capture rare modes. In contrast, MotionMap captures both common and rare mode since they are encoded in the form of local maxima.
Figure 12: We overlay predictions for each baseline on the ground truth heatmap, for each of the three input pose sequences. The encoding of these predictions is shown as red crosses. For MotionMap, we directly overlay the predicted MotionMap (with crosses for maxima) on the ground truth heatmap. We note that methods are either highly diverse but unrealistic or are less diverse but predict likely futures. In contrast, MotionMap predicts both: common and rare modes since both are explicitly encoded in the training process.